Two loudspeakers, 5.0 m apart and facing each other, play identical sounds of the same frequency. You stand halfway between them, where there is a maximum of sound intensity. Moving from this point toward one of the speakers, you encounter a minimum of sound intensity when you have moved 0.35 m . Assume the speed of sound is 340 m/s.

Part A: What is the frequency of the sound?

Express your answer with the appropriate units.

Part B: If the frequency is then increased while you remain 0.35 m from the center, what is the first frequency for which that location will be a maximum of sound intensity?

Express your answer with the appropriate units.
Answer #1

L1 = (L/2) + 0.35 = 2.5 + 0.35 = 2.85 m

L2 = 2.5 - 0.35 = 2.15 m

Path difference, phi = 2.85 - 2.15 = 0.70 m

Lambda = 2 x 0.7 = 1.4 m

Frequency, f = 340/1.4 = 242.85 hz

B)

Frequency, f = 340/0.7 = 485.71 hz

Comment in case any doubt please rate my answer ...